# Relational Graph Transformer

### v1

@@ 3.2 Transformer Network
The Transformer network in RelGT, shown in Figure 3, processes the tokenized relational entity graph using a combination of local and global attention mechanisms, following the successful design components used in modern GTs [41, 50, 30, 11].
Local module. The local attention mechanism allows each seed node to attend to its $K$ local tokens selected during tokenization, capturing the fine-grained relationships defined by the database schema. This mechanism is different from a GNN used in RDL [42] in two key aspects: self-attention is used as the message-passing scheme and the attention is all-pair, i.e., all nodes in the local K set attend to each other. This is implemented using an L layer Transformer network [47] and provides a broader structural coverage compared to a baseline GNN [42]. A practical application of this improvement can be seen in the e-commerce example introduced in Section 1, where the proposed full-attention mechanism can directly connect seemingly unrelated products by identifying relationships through shared transactions or customer behaviors. This capability enables the model to capture subtle associations, such as customers frequently purchasing unexpected combinations of items.
The local node representation $h_{\text{local}}(v_{i})$ is obtained as:
@@
where, $L$ denotes the layers, FFN and Attention are standard components in a Transformer [47], and Pool denotes the aggregation of $\{v_{j}\}_{j=1}^{K}$ and $v_{i}$ using a learnable linear combination.
Global module. The global attention mechanism enables each seed node to attend to a set of $B$ global tokens representing centroids of all nodes in the graph, conceptually and is adapted from prior works [30, 11]. These centroids are updated during training using an Exponential Moving Average (EMA) K-Means algorithm applied to seed node features in each mini-batch, providing a broader contextual view beyond the local neighborhood. The global representation is formulated as:
@@ 4 Experiments
RelGT is evaluated on the recently introduced Relational Deep Learning Benchmark (RelBench) [42]. RelBench consists of 7 datasets from diverse relational database domains, including e-commerce, clinical records, social networks, and sports, among others. These datasets are curated from their respective source domains and consist a wide range of sizes, from 1.3K to 5.4M records in the training set for the prediction tasks, with a total of 47M training records. For each dataset, multiple predictive tasks are defined, such as predicting a user’s engagement with an advertisement within the next four days or determining whether a clinical trial will achieve its primary outcome within the next year. In total, RelBench has 30 tasks across the 7 datasets, covering entity classification, entity regression, and recommendation. For our evaluation, we focus on 21 tasks on entity classification and regression  [^3] We exclude recommendation tasks in this work since they involve specific considerations, such as identifying target nodes [54] or using pair-wise learning architectures [55] and using RelGT trivially in RDL is sub-optimal..
@@ 4.1 Setup and Baselines
We implement RelGT within the RDL pipeline [42] by replacing the original GNN component, while preserving the learning mechanisms, database loaders, and task evaluators. The model has between 10-20 million parameters, and we use a learning rate of $41e-$ 4. For tasks with fewer than one million training nodes, we tune the number of layers $L\in{1,4,8}$ and use dropout rates of ${0.3,0.4,0.5}$. For tasks with more than one million training nodes, we fix the number of layers to $L=4$ due to compute budgets. For the sampling during the token preparation stage, we use $K=300$ local neighbors and set $B=4096$ as the number of tokens for global centroids. For smaller datasets (under one million training nodes), we use a batch size of 256 to ensure sufficient training steps. For larger datasets, we use a batch size of 1024. We do not perform exhaustive hyperparameter tuning; rather, our goal is to showcase the benefits of using RelGT in place of GNNs within the RDL framework. As shown in our ablation of the multi-element tokenization and global module in RelGT (Table 2), and context size (Figure 4), careful tuning may further improve performance across different tasks.
In addition to the HeteroGNN baseline used in RDL, we report results for two variants of the Heterogeneous Graph Transformer (HGT) [20] to highlight the advantages of RelGT over existing GT models. Notably, many GTs, such as GraphGPS [41], are not directly applicable to heterogeneous graphs. Therefore, we adopt HGT and an enhanced version, HGT+PE, which incorporates Laplacian positional encodings (LapPE). These positional encodings are computed on the sampled subgraphs rather than the entire graph. Additional implementation details are included in Appendix A.3.
Table 1: Test set results on the entity regression and classification tasks in RelBench. Best values are in bold. RDL: HeteroGNN baseline [42], HGT: Heterogeneous GT [20], PE: Laplacian Positional Encodings [10]. Relative gains are expressed as percentage improvement over RDL baseline.
@@
rel-f1 driver-position 4.022 4.1598 4.2358 3.9170 2.61
rel-avito ad-ctr 0.041 0.0441 0.0494 0.0345 15.85
rel-event user-attendance 0.258 0.2635 0.2562 0.2502 2.79
rel-trial study-adverse 44.473 43.3253 42.4622 43.9923 1.08
site-success 0.400 0.4374 0.4431 0.3263 18.43
rel-amazon user-ltv 14.313 15.3804 15.9296 14.2665 0.32
item-ltv 50.053 56.1384 55.6211 48.9222 2.26
@@
rel-f1 driver-dnf 0.7262 0.7142 0.7109 0.7587 4.48
driver-top3 0.7554 0.6389 0.8340 0.8352 10.56
rel-avito user-clicks 0.6590 0.6584 0.6387 0.6830 3.64
user-visits 0.6620 0.6426 0.6507 0.6678 0.88
rel-event user-repeat 0.7689 0.6717 0.6590 0.7609 -1.04
user-ignore 0.8162 0.8348 0.8161 0.8157 -0.06
rel-trial study-outcome 0.6860 0.5679 0.5691 0.6861 0.01
rel-amazon user-churn 0.7042 0.6608 0.6589 0.7039 -0.04
item-churn 0.8281 0.7824 0.7840 0.8255 -0.31
rel-stack user-engagement 0.9021 0.8898 0.8852 0.9053 0.35
user-badge 0.8986 0.8652 0.8518 0.8632 -3.94
rel-hm user-churn 0.6988 0.6773 0.6491 0.6927 -0.87
4.2 Results and Discussion
@@
RelGT improves over GNN in RDL. The experimental results in Tables 1(a) and 1(b) demonstrate that RelGT consistently matches or outperforms the standard GNN baseline used in RDL [42] across multiple datasets and tasks. We observe the largest improvements in rel-trial site-success (18.43%), rel-avito ad-ctr (15.85%), and rel-f1 driver-top3 (10.56%), while on rel-stack user-badge, RelGT performs below the RDL baseline by a margin of -3.94%. For all other tasks, RelGT consistently improves or matches the performance of the baseline GNN. We attribute the overall performance improvement to two key factors: (i) the broader structural coverage enabled by RelGT’s attention mechanisms as described in Section 3.2, and (ii) the fine-grained encodings employed in our tokenization scheme, which are further studied as follows and presented in Table 2.
@@
Subgraph GNN PE is critical in RelGT. In Table 2, we highlight the importance of several components in RelGT by conducting ablation studies. We remove one component at a time while preserving all others, and report the relative performance drop compared to the full RelGT model. Our results show that removing the subgraph GNN (PE), which encodes local subgraph structure (Section 3.1), leads to consistent performance degradation across all tasks. This component proves critical for disambiguating parent-child relationships when full-attention is applied, thanks to the random node features initialization [43, 26]. For instance, without the GNN (PE), products belonging to specific transactions (Figure 1) cannot be effectively captured, even when other encodings remain.
@@
Global module can bring gains depending on the task. In the same Table 2, our results of removing the global attention to the learnable centroids (Section 3.2) reveal task-dependent patterns that align with the findings reported in [30, 11]. For some tasks, such as rel-trial site-success, removing the attention to the centroids tokens leads to a substantial performance drop (-19.08%), indicating that the global database-wide context provides crucial information beyond the local neighborhood. However, for certain tasks such as rel-avito user-clicks, removing the global module actually improves performance (7.79% relative gain), suggesting that for some prediction targets, local information is sufficient, and the global context might introduce noise. These mixed results highlight the complementary nature of local and global information in relational graphs, with the latter being optional depending on the task.
@@
Ablation of other encodings. The remaining ablations in Table 2 reveal mixed results across different components. While removing explicit fine-grained encodings (node type, hop distance, and relative time) degrades performance on some tasks, it improves performance on others. For tasks with specific temporal dependencies (as detailed in the Appendix A.1), our current temporal encodings may inadvertently introduce noise. Similarly, for node type and hop distance encodings, their information might already be partially captured by other model components. Despite these variations, the full RelGT model still shows consistently superior results when averaged across all tasks. However, our findings suggest that RelGT’s performance could be further enhanced by careful tuning of these encoding components based on their task-specific importance. In particular, additional improvements can be achieved by incorporating more effective temporal encoding methods [7, 21, 23].
@@
rel-avito ad-ctr $-6.00$ $-1.14$ $-7.14$ $-3.43$ $-9.14$
rel-avito user-clicks $7.85$ $-15.15$ $5.01$ $5.77$ $8.37$
rel-avito user-visits $-0.35$ $-2.38$ $-0.11$ $0.39$ $-0.75$
rel-event user-ignore $-1.30$ $0.12$ $-0.11$ $0.66$ $-0.09$
rel-trial study-outcome $-2.14$ $-1.72$ $3.74$ $-0.43$ $2.48$
rel-trial site-success $-19.01$ $-9.17$ $-2.88$ $-21.49$ $-0.71$
rel-amazon user-churn $-0.64$ $-0.78$ $0.16$ $0.06$ $-2.20$
rel-hm item-sales $-9.33$ $-17.35$ $-12.69$ $0.93$ $-77.24$
Average $-3.87$ $-5.95$ $-1.75$ $-2.19$ $-9.91$
Figure 4: Left: Epoch runtime comparison of HGT [20] and HGT+PE, with Laplacian PE (see Figure 5 for all tasks). The red portion shows the additional time consumed by the precomputation of Laplacian PE against the base HGT time (blue). Right: Ablation for different $K$ values as the local context size in RelGT. Results using $K=300$ serve as the baseline (100% performance), with $K=100$ and $K=500$ runs measured as % of performance relative to $K=300$.
HGT, a GT baseline, underperforms with significant computational overhead. As shown in Tables 1(a) and 1(b), HGT [20] underperforms compared to the HeteroGNN baseline of RDL [42] across most tasks, with only two exceptions: rel-trial study-adverse and rel-event user-ignore. Notably, the integration of Laplacian eigenvectors as PEs in HGT improves performance in just 5 out of 21 tasks. Moreover, as illustrated in Figure 4, the computational overhead required for precomputing the Laplacian PEs substantially increases per-epoch runtime across various tasks. These empirical findings clearly reveal the difficulties of directly applying existing GT architectures to relational entity graphs, emphasizing the importance and need for our contributions with RelGT.
@@ 5 Conclusion
In this work, we introduce the first Graph Transformer designed specifically for relational entity graphs: the Relational Graph Transformer RelGT. It addresses key challenges faced by existing models, such as incorporating heterogeneity, temporality, and comprehensive structural modeling within a unified GT framework. RelGT represents nodes as multi-element tokens enriched with fine-grained graph context and combines local attention over sampled subgraph tokens with global attention to learnable centroids, enabling effective representation learning on relational data. Experiments on the RelBench benchmark show that RelGT consistently outperforms GNN and GT baselines across multiple tasks. Moreover, our analysis highlights the critical role of subgraph-based positional encodings as a lightweight and effective alternative to traditional graph positional encodings. This work establishes RelGT as a powerful architecture for relational deep learning and opens new avenues for advancing and scaling such architectures toward foundation models tailored for relational data.
Acknowledgments and Disclosure of Funding
@@ A.1 Benchmark Details
In this section, we include the details on the datasets and the tasks in RelBench [42] which we use for our evaluation. RelBench consists of 7 datasets from diverse relational database domains, including e-commerce, clinical records, social networks, and sports, among others. These datasets are curated from their respective source domains and consist a wide range of sizes, from 1.3K to 5.4M records in the training set for the prediction tasks, with a total of 47M training records. For each dataset, multiple predictive tasks are defined, such as predicting a user’s engagement with an advertisement within the next four days or determining whether a clinical trial will achieve its primary outcome within the next year. In total, RelBench has 30 tasks across the 7 datasets, covering entity classification, entity regression, and recommendation. For our evaluation, we focus on 21 tasks on entity classification and regression as RelGT primarily serves as a node representation learning model in RDL. We exclude recommendation tasks in this work since they involve specific considerations, such as identifying target nodes [54] or using pair-wise learning architectures [55] and using RelGT trivially in RDL is sub-optimal. We detail the dataset and task statistics in Table 3.

### v2

@@ 3.2 Transformer Network
The Transformer network in RelGT, shown in Fig. 3, processes the tokenized relational entity graph using a combination of local and global attention mechanisms, following the successful designs used in modern GTs (Rampášek et al., 2022; Wu et al., 2023; Kong et al., 2023; Dwivedi et al., 2023).
Local module. The local attention mechanism allows each seed node to attend to its $K$ local tokens selected during tokenization, capturing the fine-grained relationships defined by the database schema. This mechanism is different from a GNN used in RDL (Robinson et al., 2024) in two key aspects: self-attention is used as the message-passing scheme and the attention is all-pair, i.e., all nodes in the local K set attend to each other. This is implemented using an L layer Transformer (Vaswani et al., 2017) and provides a broader structural coverage compared to a baseline GNN (Robinson et al., 2024). A practical application of this improvement can be seen in the e-commerce example introduced in Sec. 1, where the proposed full-attention mechanism can directly connect seemingly unrelated products by identifying relationships through shared transactions or customer behaviors. This capability enables the model to capture subtle associations, such as customers frequently purchasing unexpected combinations of items. The local node representation $h_{\text{local}}(v_{i})$ is obtained as:
@@
where, $L$ denotes the layers, FFN and Attention are standard components in a Transformer (Vaswani et al., 2017), and Pool denotes the aggregation of $\{v_{j}\}_{j=1}^{K}$ and $v_{i}$ using a learnable linear combination.
Global module. The global attention mechanism enables each seed node to attend to a set of $B$ global tokens representing centroids of all nodes in the graph, conceptually and is adapted from prior works (Kong et al., 2023; Dwivedi et al., 2023). These centroids are updated during training using an Exponential Moving Average (EMA) K-Means algorithm applied to seed node features in each mini-batch, providing a broader contextual view beyond the local neighborhood. The global representation is formulated as:
@@ 4 Experiments
RelGT is evaluated on the recently introduced RDL Benchmark (RelBench) (Robinson et al., 2024). RelBench consists of 7 datasets from diverse relational database domains, including e-commerce, clinical records, social networks, and sports, among others. These datasets are curated from their respective source domains and consist a wide range of sizes, from 1.3K to 5.4M records in the training set for the prediction tasks, with a total of 47M training records. For each dataset, multiple predictive tasks are defined, such as predicting a user’s engagement with an advertisement within the next four days or determining whether a clinical trial will achieve its primary outcome within the next year. In total, RelBench has 30 tasks across the 7 datasets, covering entity classification, entity regression, and recommendation. For our evaluation, we focus on 21 tasks on entity classification and regression  [^3] We exclude recommendation tasks in this work since they involve specific considerations, such as identifying target nodes (You et al., 2021) or using pair-wise learning architectures (Yuan et al., 2024). Details in Sec. A.1.
@@ 4.1 Setup and Baselines
We implement RelGT within the RDL pipeline (Robinson et al., 2024) by replacing the original GNN component, while preserving the learning mechanisms, database loaders, and task evaluators. The model has between 10-20M parameters, and we use a learning rate of 1e41e-4. For tasks with fewer than 1M training nodes, we tune the number of layers $L\in{1,4,8}$ and use dropout rates of 0.3,0.4,0.5{0.3,0.4,0.5}. For tasks with more than 1M training nodes, we fix the number of layers to $L=4$ due to compute budgets. For the sampling during the token preparation stage, we use $K=300$ local neighbors and set $B=4096$ as the number of tokens for global centroids. For smaller datasets (under one million training nodes), we use a batch size of 256 to ensure sufficient training steps. For larger datasets, we use a batch size of 1024. We do not perform exhaustive hyperparameter tuning; rather, our goal is to showcase the benefits of using RelGT in place of GNNs within the RDL framework. As shown in our ablation of the multi-element tokenization and global module in RelGT (Tab.  2), and context size (Fig. 4), careful tuning may further improve performance across different tasks.
In addition to the HeteroGNN baseline used in RDL, we report results for two variants of the Heterogeneous Graph Transformer (HGT) (Hu et al., 2020) to highlight the advantages of RelGT over existing GT models. Notably, many GTs, such as GraphGPS (Rampášek et al., 2022), are not directly applicable to heterogeneous graphs. Therefore, we adopt HGT and an enhanced version, HGT+PE, which incorporates Laplacian PE. These positional encodings are computed on the sampled subgraphs rather than the entire graph. Additional details are included in Appendix A.5.
4.2 Results and Discussion
Table 1: Test set results on the entity regression and classification tasks in RelBench. Best values are in bold. RDL: HeteroGNN baseline (Robinson et al., 2024), HGT: Heterogeneous GT (Hu et al., 2020), PE: Laplacian PE. Relative gains are expressed as percentage improvement over RDL baseline.
@@ 4.2 Results and Discussion
rel-f1 driver-position 4.022 4.2263 4.3921 3.9170 2.61
rel-avito ad-ctr 0.041 0.0462 0.0483 0.0345 15.85
rel-event user-attendance 0.258 0.2635 0.2611 0.2502 2.79
rel-trial study-adverse 44.473 45.1692 42.6484 43.9923 1.08
site-success 0.400 0.4428 0.4396 0.3263 18.43
rel-amazon user-ltv 14.313 15.4120 15.8643 14.2665 0.32
item-ltv 50.053 55.8683 55.8493 48.9222 2.26
@@
rel-f1 driver-dnf 0.7262 0.7077 0.7117 0.7587 4.48
driver-top3 0.7554 0.7075 0.7627 0.8352 10.56
rel-avito user-clicks 0.6590 0.6376 0.6457 0.6830 3.64
user-visits 0.6620 0.6432 0.6495 0.6678 0.88
rel-event user-repeat 0.7689 0.6496 0.6536 0.7609 -1.04
user-ignore 0.8162 0.8247 0.8161 0.8157 -0.06
rel-trial study-outcome 0.6860 0.5837 0.5921 0.6861 0.01
rel-amazon user-churn 0.7042 0.6643 0.6619 0.7039 -0.04
item-churn 0.8281 0.7797 0.7803 0.8255 -0.31
rel-stack user-engagement 0.9021 0.8847 0.8817 0.9053 0.35
user-badge 0.8986 0.8608 0.8566 0.8632 -3.94
rel-hm user-churn 0.6988 0.6695 0.6569 0.6927 -0.87
@@
RelGT improves over GNN in RDL. The experimental results in Tables 1(a) and 1(b) demonstrate that RelGT consistently matches or outperforms the standard GNN baseline used in RDL (Robinson et al., 2024) across multiple datasets and tasks. Using a ±1% threshold to assess comparable performance, RelGT shows: (i) clear improvements (more than a 1% relative gain) on 10 tasks, (ii) comparable results (within ±1%) on 9 tasks, and (iii) competitive but lower performance (more than a 1% relative loss) on 2 tasks. We observe the largest improvements in rel-trial site-success (18.43%), rel-avito ad-ctr (15.85%), and rel-f1 driver-top3 (10.56%), while on rel-stack user-badge, RelGT performs below the RDL baseline by a margin of -3.94%. For all other tasks, RelGT consistently improves or matches the performance of the baseline GNN. We attribute the overall performance improvement to two key factors: (i) the broader structural coverage enabled by RelGT’s attention mechanisms as described in Section 3.2, and (ii) the fine-grained encodings employed in our tokenization scheme, which are further studied as follows and presented in Table 2.
@@
Subgraph GNN PE is critical in RelGT. In Table 2, we highlight the importance of several components in RelGT by conducting ablation studies. We remove one component at a time while preserving all others, and report the relative performance drop compared to the full RelGT model. Our results show that removing the subgraph GNN (PE), which encodes local subgraph structure (Section 3.1), leads to consistent performance degradation across all tasks. This component proves critical for disambiguating parent-child relationships when full-attention is applied, thanks to the random node features initialization (Sato et al., 2021; Kanatsoulis et al., 2025). For instance, without the GNN (PE), products belonging to specific transactions (Figure 1) cannot be effectively captured, even when other encodings remain.
@@
Global module can bring gains depending on the task. In the same Table 2, our results of removing the global attention to the learnable centroids (Section 3.2) reveal task-dependent patterns that align with the findings reported in (Kong et al., 2023; Dwivedi et al., 2023). For some tasks, such as rel-trial site-success, removing the attention to the centroids tokens leads to a substantial performance drop (-19.08%), indicating that the global database-wide context provides crucial information beyond the local neighborhood. However, for certain tasks such as rel-avito user-clicks, removing the global module actually improves performance (7.79% relative gain), suggesting that for some prediction targets, local information is sufficient, and the global context might introduce noise. These mixed results highlight the complementary nature of local and global information in relational graphs, with the latter being optional depending on the task.
@@
Ablation of other encodings. The remaining ablations in Table 2 reveal mixed results across different components. While removing explicit fine-grained encodings (node type, hop distance, and relative time) degrades performance on some tasks, it improves performance on others. For tasks with specific temporal dependencies (detailed in Sec. A.1), our current temporal encodings may inadvertently introduce noise. Similarly, for node type and hop distance encodings, their information might already be partially captured by other model components. Despite these variations, the full RelGT model still shows consistently superior results when averaged across all tasks. However, our findings suggest that RelGT’s scores could be further enhanced by careful tuning of these encoding components based on their task-specific importance. In particular, additional gains can be achieved by incorporating more effective temporal encodings (Clauset & Eagle, 2012; Huang et al., 2024; Jiang & Pu, 2023).
@@
rel-avito ad-ctr 6.00-6.00 1.14-1.14 7.14-7.14 3.43-3.43 9.14-9.14
rel-avito user-clicks $7.85$ 15.15-15.15 $5.01$ $5.77$ $8.37$
rel-avito user-visits 0.35-0.35 2.38-2.38 0.11-0.11 $0.39$ 0.75-0.75
rel-event user-ignore 1.30-1.30 $0.12$ 0.11-0.11 $0.66$ 0.09-0.09
rel-trial study-outcome 2.14-2.14 1.72-1.72 $3.74$ 0.43-0.43 $2.48$
rel-trial site-success 19.01-19.01 9.17-9.17 2.88-2.88 21.49-21.49 0.71-0.71
rel-amazon user-churn 0.64-0.64 0.78-0.78 $0.16$ $0.06$ 2.20-2.20
rel-hm item-sales 9.33-9.33 17.35-17.35 12.69-12.69 $0.93$ 77.24-77.24
Average 3.87-3.87 5.95-5.95 1.75-1.75 2.19-2.19 9.91-9.91
Figure 4: Left: Epoch runtime comparison of HGT (Hu et al., 2020) and HGT+PE, with Laplacian PE (see Figure 5 for all tasks). The red portion shows the additional time consumed by the precomputation of Laplacian PE against the base HGT time (blue). Right: Ablation for different $K$ values as the local context size in RelGT. Results using $K=300$ serve as the baseline (100% performance), with $K=100$ and $K=500$ runs measured as % of performance relative to $K=300$.
HGT, a GT baseline, underperforms with significant overhead. As shown in Tables 1(a) and 1(b), HGT (Hu et al., 2020) underperforms compared to the HeteroGNN baseline of RDL (Robinson et al., 2024) across most tasks, with only two exceptions: rel-trial study-adverse and rel-event user-ignore. Notably, the integration of Laplacian PEs in HGT improves performance in 8 (of 21) tasks. Moreover, as illustrated in Figure 4, the computational overhead required for precomputing the Laplacian PEs substantially increases per-epoch runtime across various tasks. These empirical findings clearly reveal the difficulties of directly applying existing GT architectures to relational entity graphs, emphasizing the importance and need for our contributions with RelGT.
@@ 5 Conclusion
In this work, we introduce the first Graph Transformer designed specifically for relational entity graphs: the Relational Graph Transformer. It addresses key challenges faced by existing models, such as incorporating heterogeneity, temporality, and comprehensive structural modeling within a unified GT framework. RelGT represents nodes as multi-element tokens enriched with fine-grained graph context and combines local attention over sampled subgraph tokens with global attention to learnable centroids, enabling effective representation learning on relational data. Experiments on the RelBench benchmark show that RelGT consistently outperforms GNN and GT baselines across multiple tasks. Moreover, our analysis highlights the critical role of subgraph-based positional encodings as a lightweight and effective alternative to traditional graph positional encodings. This work establishes RelGT as a powerful architecture for relational deep learning and opens new avenues for advancing and scaling such architectures toward foundation models tailored for relational data.
Acknowledgments
@@ A.1 Benchmark Details
In this section, we include the details on the datasets and the tasks in RelBench (Robinson et al., 2024) which we use for our evaluation. RelBench consists of 7 datasets from diverse relational database domains, including e-commerce, clinical records, social networks, and sports, among others. These datasets are curated from their respective source domains and consist a wide range of sizes, from 1.3K to 5.4M records in the training set for the prediction tasks, with a total of 47M training records. For each dataset, multiple predictive tasks are defined, such as predicting a user’s engagement with an advertisement within the next four days or determining whether a clinical trial will achieve its primary outcome within the next year. In total, RelBench has 30 tasks across the 7 datasets, covering entity classification, entity regression, and recommendation. For our evaluation, we focus on 21 tasks on entity classification and regression as RelGT primarily serves as a node representation learning model in RDL. We exclude recommendation tasks in this work since they involve specific considerations, such as identifying target nodes (You et al., 2021) or using pair-wise learning architectures (Yuan et al., 2024) and using RelGT trivially in RDL is sub-optimal. We detail the dataset and task statistics in Table 3.